OpenStudy (anonymous):

Given one endpoint C (-5,-2) And the midpoint M (-1,2) ... Find D . The other endpoint . Helpp please

OpenStudy (anonymous):

@Calcmathlete any ideas?

OpenStudy (matheducatormcg):

graph C, then graph M. find the slope pattern from C to M and then continue the pattern one more point.

OpenStudy (campbell_st):

same as before for x for y -1 = ( -5 +x)/2 2 = (-2 + y)/2 -2 = -5 + x 4 = -2 + y solve for x and y add 5 to both sides add 2 to both sides x = 3 y = 6 ordered pair is (3, 6)